White Wedding Corporation will pay a $2.40 per share dividend next year. The company pledges to increase its dividend by 7.10 percent per year indefinitely.

Required:

If you require an 10.2 percent return on your investment, how much will you pay for the company's stock today? (Do not include the dollar sign ($). Round your answer to 2 decimal places. (e.g., 32.16))

  Current share price $   
Answer #1

Using the constant growth model, we find the price of the stock today is:

P0 = D1 / (R − g) = $2.40 / (0.102 −0.0710) = $77.42